# CosmicWatch: The Desktop Muon Detector (v3X)

### v1

@@ Abstract
Figure 1: An Array of the CosmicWatch Desktop Muon Detector (v3X).
@@ 1.1 CosmicWatch
CosmicWatch [1, 2, 3] is an University of Delaware (UD)-lead outreach program aimed at introducing students to particle, nuclear, and astro-particle physics by providing educational material and support for building particle detectors and making physics measurements.
The original program, initiated at the Massachusetts Institute of Technology (MIT) and Polish National Center for Nuclear Research (NCBJ) in 2017, has substantially grown and changed, resulting in novel uses for the detectors, including calibration instruments for neutrino detectors [4], spacecraft instrumentation, education, and outreach [5][6][7][8]. It is now substantially used as a versatile platform for teaching high-school, undergraduates, and graduate students practical skills such as soldering, micro-controller unit (MCU) programming, introducing fundamental physics concepts, and giving hands-on experience with particle detection techniques.
@@ 1.2 The Desktop Muon Detector (v3X)
Over the past year, there have been active efforts to develop a new version of the detector (v3X), shown in figure 1, with significant improvements over the previous model, v1 [3] and v2 [1, 2]. These improvements expand the physics capabilities of the detectors, enhance the performance, and add new functionalities, while keeping the total cost of a detector around \$100. A comprehensive list comparing the performance between the two generations of detectors is shown in table 1. All material pertaining to the v3X detector can be found in the GitHub repository located here:
@@ 15 Hz Maximum event rate 700 Hz
Yes (w/o SD card) OLED readout Yes
@@
Furthermore, the detector’s modular architecture readily accommodates future upgrades. With minimal modifications it can be reconfigured as a gamma-ray spectrometer or linked to an expansion module: including GPS receivers, real-time clocks, magnetometers, WiFi and Bluetooth communication, and environmental sensors. This flexibility lays the groundwork for a planned citizen-science network, enabling volunteers to map Earth’s ionizing radiation and, ultimately, to reconstruct cosmic-ray air showers in distributed arrays of detectors.
@@ 1.2.1 Principle of operation
The SiPM signal is amplified and sent to a comparator and a peak detector circuit. The comparator triggers data acquisition by providing a timestamp, while the peak detector holds a constant fraction of the pulse peak for accurate amplitude measurement by the Raspberry Pi Pico (RPi) MCU [11]. The resulting ADC value is proportional to the number of incident photons. Additionally, the SiPM signal is available through a BNC connector for external readout with nanosecond timing resolution (see Sec.6.4).
@@
To suppress the radiogenic background signals and focus on cosmic-ray muons, the v3X detector supports coincidence detection using multiple units. In a two-detector configuration, each unit independently processes events as previously described; however, a coincidence event, defined as a trigger in both detectors within a predefined time window, is flagged when both comparators fire nearly simultaneously. This temporal correlation effectively reject uncorrelated background events, such as environmental gamma and beta radiation, which are unlikely to occur in both spatially separated detectors at the same time. Coincidence detection enables the identification of through-going cosmic-ray muons and forms the basis for more advanced applications, including multi-detector muon tracking geometries and energy deposition studies (see Sec. 6.1), and depth and altitude measurements (see Sec. 6.3).
@@
Once triggered, the RPi records the event count, timestamp, 12-bit pulse amplitude, SiPM pulse height, temperature, barometric pressure, linear acceleration, angular velocity, cumulative dead time, and whether or not the event was seen by a coincident detector. The master and coincidence count rates are shown on an OLED display, while every event’s data is recorded to a microSD card and streamed in real time via USB serial port. If data is saved directly to a computer using the provided software, an additional time and date stamp are provided through Network Time Protocol (NTP). This enables global synchronized real-time information and removes the internal drift of the MCU oscillator. All data is saved as easy to read, tab-delimited .txt file, with each row corresponding to an individual triggering event. Once an event has been fully recorded, a metal-oxide-semiconductor field-effect transistor (MOSFET) shunts the peak detector charging capacitor to ground, initiating the search for subsequent pulses.
@@
Although users do not have direct access to the firmware, when a microSD card is inserted, a file named configure.txt is created, which contains variables that can be changed to provide some level of flexibility. This includes options to change the detector trigger threshold, rename the detector, turn off the OLED screen and external LED lights, input the wifi address and password (if built using a RPi W), and enable/disable the buzzer. Once modified, when the detector reboots, it will account for these modifications.
@@
By connecting two v3X detectors using a standard CAT5/6 cable via the coincidence port (see figure 1), users can enable coincidence detection, enabling accurate identification of cosmic-ray muons and selection of the incoming particle’s direction.
@@ 2.1 Scintillator and single photon sensor
Scintillation photons are either internally reflected at the scintillator–air interface or, if they escape, redirected by a surrounding layer of reflective aluminum foil. A portion of these photons reach a photosensor– OnSemiconductor MicroFC 60035 C-Series SiPM–which measures $\times7 mm$, with a $\times6 mm$ active area [15]. The SiPM is mounted on a custom 2-layer PCB and operated at 30 V reverse bias. This corresponds to an overvoltage (voltage above the breakdown voltage) of 5.5 V, achieving a photon detection efficiency of about 43% at 420 nm.
@@ 2.2.3 Digital electronics
The system supports USB serial communication (at a baud rate of 115,200 bps), enabling real-time monitoring and firmware updates. When data is captured via USB serial port on a computer using the provided import_data.py software, each event is timestamped in real time by the host computer. These timestamps typically synchronize to the NTP, ensuring accuracy within hundreds of microseconds to a few milliseconds.
@@
Collectively, these digital electronics create a flexible, low-latency platform suited for detector control, efficient data acquisition, and straightforward integration with external peripherals.
@@ 5.2.1 OLED readout
The OLED display is ideal for users who only need overall event counts and average event rates, without detailed per-event data. It reports the total run time, number of triggered events, and the average count rate (accounting for dead time). If the detector boots into coincidence mode, two additional rows indicating the number of coincident events and coincidence average count rate are also reported. Additionally, the OLED display reports the microSD card file name (if present), and the detector name (modifiable in the configure.txt file located on the microSD card). The OLED can also be turned on/off using the configure.txt file.
@@ 6.1 A full-sky cosmic-ray muon measurement
Figure 10: Left: the SiPM pulse-height spectra recorded by a set of detector in configuration (d). The coincident events begin dominating at SiPM pulse-heights greater than approximately 18 mV. Right: The measured detection rate of coincident and non-coincident events.
The SiPM pulse-height spectrum (figure 10, left), converted from ADC counts using the calibration curve in figure 6, exhibits the characteristic Landau-shaped “minimum ionizing particle (MIP) bump” produced by through-going muons in a thin absorber. The Landau energy-loss distribution can be approximated by the analytic Moyal distribution [16, 17]:
@@
where $\mu$ denotes the most probable value and $c$ the width parameter. A fit over the range $13\,\mathrm{mV}$ to $40\,\mathrm{mV}$ yields $\mu\approx 21\,\mathrm{mV}$, which corresponds to the $\sim 2\,\mathrm{MeV}$ energy loss of a minimum-ionizing muon traversing the $\approx 1\,\mathrm{cm}$-thick plastic scintillator [18], and $c\approx 5\,\mathrm{mV}$. This is only approximate, as the average trajectory of the muons through the scintillator will be slightly larger than $1\,\mathrm{cm}$. Detector saturation appears near $180\,\mathrm{mV}$, as denoted by the overflow bin. The relatively flat tail below $10\,\mathrm{mV}$ in the coincident-event distribution is attributed to corner-clipping muons, which graze the scintillator edge and deposit reduced energy.
This demonstrates the ability to select muons using coincidence mode and shows how the SiPM pulse amplitude can be used to further discriminate between muons and radiogenic backgrounds based on energy deposition. The ratio plotted in the left subplot of figure 10 indicates that above approximately $18\,\mathrm{mV}$, cosmic-ray muons dominate the energy deposition spectrum.
The precise location of the MIP peak depends on the precise value of the breakdown voltage as well as the quality of the detector assembly: the efficiency of the optical coupling between the SiPM and scintillator, the scintillator surface finish, the reflectivity of any wrapping, and other factors that influence photon collection. It also shifts with the SiPM bias voltage, which is set by precision (0.1 %) resistors and directly determines the gain. Nevertheless, because the MIP peak occurs well above the trigger threshold ($\approx$4 mV), even substantial variations in collection efficiency have minimal effect on muon detection efficiency; only corner-clipping events register near threshold. Overall count rates, however, are expected to show more significant detector to detector variations due to the steep rise of the non-coincident background near threshold.
Figure 10 (right) displays the event rate versus time for the same dataset, comparing non-coincident and coincident counts. Each point represents a 240 s interval, with the rate calculated as the number of events divided by the effective live time (bin duration minus total dead time). These rate measurements enable searches for correlations with atmospheric variables; for e.g., barometric pressure changes (figure 11 (left)) or transient phenomena such as solar flares. The detector also logs temperature (figure 11 [right]), allowing corrections for temperature-dependent gain shifts, although it currently isn’t implemented.
@@ 6.2 Cosmic-ray muon angular dependence
This measurement illustrates the cosmic-ray muon angular dependence. Two detectors were operated in coincidence mode and positioned side-by-side (configuration (a) in figure 7), with their scintillators separated by 50 mm inside identical aluminum enclosures. The detector angle relative to the zenith, $\theta$, was set by mounting both units on a 40 cm rectangular bar and leaning the bar against a vertical wall at a known height (see figure 12, right). Figure 12 (left) presents the measured coincidence rate with statistical error bars, where each data point corresponds to roughly 24 hours of data acquisition.
@@ 6.4 Cosmic-ray muon velocity
The TOF measurement measurement shown in figure 14 recorded 92 events over the course of several weeks with a pulse separation of less than 10 ns. A Gaussian fit to the resulting velocity distribution yielded a mean muon velocity of $0.92\pm 0.09\,c$, where $\mu$ and $\sigma$ correspond to the mean and standard deviation, respectively. The expected number of accidental coincidence events, calculated using equation 6.2, was found to be less than one over the duration of the measurement.
@@ 8 Acknowledgments
The authors gratefully acknowledge Amanda Meng, Liam Roth, Maeve Owens, Musarate Shams, Michael Manfre, Christina Love, Rafael Hurtado Carrillo, and Andres Felipe Vargas for their assistance with the measurements presented in this work. We also thank our colleagues at MIT, WIPAC, and Harvard for their support and for helping to integrate this effort into a high-school and undergraduate research initiative.

### v2

@@ Abstract
Figure 1: An array of the CosmicWatch Desktop Muon Detectors v3X. The front and rear faceplates provide access to the SiPM output, microSD card slot, microUSB connection, and coincidence ports. Additional elements include an OLED display, reset button, trigger LED, and coincidence LED.
@@ 1.1 CosmicWatch
CosmicWatch [^3] https://sites.udel.edu/cosmicwatch/ [1, 2, 3] is an University of Delaware (UD)-lead outreach program aimed at introducing students to particle, nuclear, and astro-particle physics by providing educational material and support for building particle detectors and making physics measurements. The original program, initiated at the Massachusetts Institute of Technology (MIT) and Polish National Center for Nuclear Research (NCBJ) in 2017, has substantially grown and changed, resulting in novel uses for the detectors, including calibration instruments for neutrino detectors [4], spacecraft instrumentation, education, and outreach [5][6][7][8]. It is also substantially used as a versatile platform for teaching high-school, undergraduates, and graduate students practical skills such as soldering, micro-controller unit (MCU) programming, introducing fundamental physics concepts, and giving hands-on experience with particle detection techniques.
@@ 1.2 The Desktop Muon Detector (v3X)
This paper presents the v3X Desktop Muon Detector, shown in figure 1, with significant improvements over the previous models, v1 [3] and v2 [1, 2]. These improvements expand the physics capabilities of the detectors, enhance the performance, and add new functionalities, while keeping the total cost of a detector around \$100. A comprehensive list comparing the performance between the two generations of detectors is shown in table 1. All material pertaining to the v3X detector can be found in the GitHub repository:
@@ 15 Hz Maximum event rate 700 Hz
Yes (w/o SD card) OLED display Yes
@@
Furthermore, the detector’s modular architecture readily accommodates future upgrades. With minimal modifications it can be reconfigured as a gamma‐ray spectrometer or linked to an expansion module: including GPS receivers, real‐time clocks, magnetometers, WiFi and Bluetooth communication, and environmental sensors. This flexibility lays the groundwork for a planned citizen‐science network, enabling volunteers to map Earth’s ionizing radiation and, ultimately, to reconstruct cosmic‐ray air showers in distributed arrays of detectors.
@@ 1.2.1 Principle of operation
The SiPM signal is amplified and sent to a comparator and a peak detector circuit. The comparator triggers data acquisition by providing a timestamp, while the peak detector holds a constant fraction of the pulse peak for accurate amplitude measurement by the Raspberry Pi Pico (RPi) MCU [11]. The resulting ADC value is proportional to the number of incident photons. Additionally, the SiPM signal is available through a BNC connector for external readout with nanosecond timing resolution (see Sec.6.4). A block diagram of the main board is illustration in Fig. 2.
Figure 2: Block diagram of the main board. The analog section shows the amplification, peak detection, as well as the comparator. The digital section illustrates the communication between peripherals as well as monitors and controls.
@@
To suppress the radiogenic background signals and focus on cosmic-ray muons, the v3X detector supports coincidence detection using multiple units. In a two-detector configuration, each unit independently processes events as previously described; however, a coincidence event, defined as a trigger in both detectors within a predefined time window, is flagged when both comparators fire nearly simultaneously. This temporal correlation effectively reject uncorrelated background events, such as environmental gamma and beta radiation, which are unlikely to occur in both spatially separated detectors at the same time. Coincidence detection enables the identification of through-going cosmic-ray muons and forms the basis for more advanced applications, including muon tracking geometries and energy deposition studies (see Sec. 6.1), and altitude measurements (see Sec. 6.3).
@@
Once triggered, the RPi records the event count, timestamp, 12-bit pulse amplitude, SiPM pulse height, temperature, barometric pressure, linear acceleration, angular velocity, cumulative dead time, and whether or not the event was seen by a coincident detector. The master and coincidence count rates are shown on an OLED display, while every event’s data is recorded to a microSD card and streamed in real time via USB port. If data is saved directly to a computer using the provided software, an additional time and date stamp are provided through Network Time Protocol (NTP). This enables global synchronized real-time information and removes the internal drift of the MCU oscillator. All data is saved as easy to read, tab-delimited .txt file, with each row corresponding to an individual triggering event. Once an event has been fully recorded, a metal-oxide-semiconductor field-effect transistor (MOSFET) shunts the peak detector charging capacitor to ground, initiating the search for subsequent pulses.
@@
Although users do not have direct access to the firmware, when a microSD card is inserted, a file named configure.txt is created, which contains variables that can be changed to provide some level of flexibility. This includes options to change the detector trigger threshold, rename the detector, turn off the OLED screen and external LED lights, and enable/disable the buzzer.
@@
By connecting two v3X detectors using a standard CAT5/6 cable via the coincidence port (see figure 1), users can enable coincidence detection, enabling accurate identification of cosmic-ray muons and selection of the incoming particle’s direction. The muon rate at the sea-level is approximately $\mathrm{cm^{-2}min^{-1}} [1$2], which corresponds to 0.42 Hz given the 25 cm2 scintillator employed in the v3X detector.
@@ 2.1 Scintillator and single photon sensor
Scintillation photons are either internally reflected at the scintillator–air interface or, if they escape, redirected by a surrounding layer of reflective aluminum foil. A portion of these photons reach a photosensor, an OnSemiconductor MicroFC (60035 C-Series) SiPM, which measures $\times7 mm$, with a $\times6 mm$ active area [16]. The SiPM is mounted on a custom 2-layer PCB and operated at 30 V reverse bias. This corresponds to an overvoltage (voltage above the breakdown voltage) of 5.5 V, achieving a photon detection efficiency of about 43% at 420 nm.
@@ 2.2.3 Digital electronics
The system supports USB serial communication (at a baud rate of 115,200 bps), enabling real-time monitoring and firmware updates. When data is captured via USB port on a computer using the provided import_data.py software, each event is timestamped in real time by the host computer. These timestamps typically synchronize to the NTP, ensuring accuracy within hundreds of microseconds to a few milliseconds.
@@
Collectively, these digital electronics create a flexible, low-latency platform suited for detector control, efficient data acquisition, and straightforward integration with external peripherals. These detectors are designed to operate reliably within a temperature range of approximately –20 °C to +85 °C.
@@ 4 Firmware Features and Functionality
Figure 8: Logic diagram of the CosmicWatch data acquisition chain. Signals from the SiPM coupled to the scintillator are amplified and passed through a peak detector. The trigger logic compares the signal to a threshold, records coincidences, and manages reset. Events are timestamped and written to the SD card buffer or printed to the serial port.
@@ 5.2.1 OLED readout
The OLED display is ideal for users who only need overall event counts and average event rates, without detailed per‑event data. It reports the total run time, number of triggered events, and the average count rate (accounting for dead time). If the detector boots into coincidence mode, two additional rows indicating the number of coincident events and coincidence average count rate are also reported. Additionally, the OLED display reports the microSD card file name (if present), and the detector name (modifiable in the configure.txt file located on the microSD card). The OLED can also be turned on/off using the configure.txt file.
@@ 6.1 A full-sky cosmic-ray muon measurement
Figure 12: Left: the SiPM pulse‐height spectra recorded by a set of detector in configuration (d). The coincident events begin dominating at SiPM pulse-heights greater than approximately 18 mV. Right: The measured detection rate of coincident and non-coincident events.
The SiPM pulse‐height spectrum (figure 12, left), converted from ADC counts using the calibration curve in figure 7, exhibits the characteristic Landau‐shaped “minimum ionizing particle (MIP) bump” produced by through‐going muons in a thin absorber. The Landau energy‐loss distribution can be approximated by the analytic Moyal distribution [17, 18]:
@@
where $\mu$ denotes the most probable value and $c$ the width parameter. A fit over the range $13\,\mathrm{mV}$ to $40\,\mathrm{mV}$ yields $\mu\approx 21\,\mathrm{mV}$, which corresponds to the $\sim 2\,\mathrm{MeV}$ energy loss of a minimum‐ionizing muon traversing the $\approx 1\,\mathrm{cm}$-thick plastic scintillator [19], and $c\approx 5\,\mathrm{mV}$. This is only approximate, as the average trajectory of the muons through the scintillator will be slightly larger than $1\,\mathrm{cm}$. Detector saturation appears near $180\,\mathrm{mV}$, as denoted by the overflow bin. The relatively flat tail below $10\,\mathrm{mV}$ in the coincident‐event distribution is attributed to corner‐clipping muons, which graze the scintillator edge and deposit reduced energy.
This demonstrates the ability to select muons using coincidence mode and shows how the SiPM pulse amplitude can be used to further discriminate between muons and radiogenic backgrounds based on energy deposition. The ratio plotted in the left subplot of figure 12 indicates that above approximately $18\,\mathrm{mV}$, cosmic‑ray muons dominate the energy deposition spectrum.
The precise location of the MIP peak depends on the precise value of the breakdown voltage as well as the quality of the detector assembly: the efficiency of the optical coupling between the SiPM and scintillator, the scintillator surface finish, the reflectivity of any wrapping, and other factors that influence photon collection. It also shifts with the SiPM bias voltage, which is set by precision (0.1 %) resistors and directly determines the gain. Nevertheless, because the MIP peak occurs well above the trigger threshold ($\approx$4 mV), even substantial variations in collection efficiency have minimal effect on muon detection efficiency; only corner‑clipping events register near threshold. Overall count rates, however, are expected to show more significant detector to detector variations due to the steep rise of the non‑coincident background near threshold.
Figure 12 (right) displays the event rate versus time for the same dataset, comparing non‑coincident and coincident counts. Each point represents a 240 s interval, with the rate calculated as the number of events divided by the effective live time (bin duration minus total dead time). These rate measurements enable searches for correlations with atmospheric variables; for e.g., barometric pressure changes (figure 13 (left)) or transient phenomena such as solar flares. The detector also logs temperature (figure 13 [right]), allowing corrections for temperature‑dependent gain shifts, although it currently isn’t implemented.
@@ 6.2 Cosmic-ray muon angular dependence
This measurement illustrates the cosmic‐ray muon angular dependence. Two detectors were operated in coincidence mode and positioned side‐by‐side (configuration (a) in figure 9), with their scintillators separated by 50 mm inside identical aluminum enclosures. The detector angle relative to the zenith, $\theta$, was set by mounting both units on a 40 cm rectangular bar and leaning the bar against a vertical wall at a known height (see figure 14, right). Figure 14 (left) presents the measured coincidence rate with statistical error bars, where each data point corresponds to roughly 24 hours of data acquisition.
@@ 6.4 Cosmic-ray muon velocity
The TOF measurement shown in figure 16 recorded 92 events over the course of several weeks with a pulse separation of less than 10 ns. A Gaussian fit to the resulting velocity distribution yielded a mean muon velocity of $0.92\pm 0.09\,c$, where $\mu$ and $\sigma$ correspond to the mean and standard deviation, respectively. The expected number of accidental coincidence events, calculated using equation 6.2, was found to be less than one over the duration of the measurement.
@@ 8 Acknowledgments
The authors gratefully acknowledge Amanda Meng, Liam Roth, Maeve Owens, Musarate Shams, Michael Manfre, Christina Love, Rafael Hurtado Carrillo, and Andres Felipe Vargas for their assistance with the measurements presented in this work. We also thank our colleagues at MIT, WIPAC, and Harvard for their support and for helping to integrate this effort into a high‑school and undergraduate research initiative.